# A Note on the Core of 2-Matching Games

### v1

@@ 2 Separating over the Core
Here, $\mathcal{C}$ stands for the set of cycles $C\subseteq E$ in $G$ with $b_{i}=2$ for all $i\in V(C)$, and $\mathcal{P}$ stands for the set of paths $P\subseteq E$ with $b_{i}=2$ for all inner vertices on $P$. We shortened $p(V(C))$ and $p(V(P))$ to $p(C)$ and $p(P)$, respectively. With this observation, separating over the core for a given vector $p$ reduces to checking whether $p(N)=\nu(N)$, which can be done in polynomial time [^1] A maximum-weight $b$-matching in a graph $G$ can be computed in polynomial time by [11]., and to separating over the set of constraints for cycles, and the set of constraints for paths.
@@
$i_{0}j^{(1)}$ for $\text{for }i_{0}j\in E$ $\text{with weight }\overline{w}_{i_{0}j}:=p_{i_{0}}+p_{j}/2-w_{i_{0}j},$ (5)
$i^{(r-1)}j^{(r)}$ for $\text{for }ij\in E$ $\text{with weight }\overline{w}_{ij}:=(p_{i}+p_{j})/2-w_{ij},$ (6)
$i^{(k-1)}j_{0}$ for $\text{for }ij_{0}\in E$ $\text{with weight }\overline{w}_{ij_{0}}:=p_{i}/2+p_{j_{0}}-w_{ij_{0}}.$ (7)
@@
$\begin{split}&p_{i}\geq 0,\quad\text{for all }i\in N,\\ &p_{i}+p_{j}\geq w_{ij},\quad\text{for all }ij\in E.\end{split}$ (9)
@@
$\begin{split}&w(C)=w_{e_{1}}+\cdots+w_{e_{k}}+w_{st}=w(P)+0=w(P),\\ &p^{\prime}(C)=p^{\prime}_{e_{1}}+\cdots+p^{\prime}_{e_{k}}+p^{\prime}_{st}=p(P),\end{split}$ (10)
@@
$\begin{split}&w(P)=w(C)-w_{st}=w(C)-0=w(C),\\ &p(P)=p^{\prime}(C),\end{split}$ (11)
@@ 3 Compact Extended Formulation
$w(C)/p^{\prime}(C)>1\iff 0>\sum_{ij\in C}c_{ij}=\sum_{ij\in C}\left(p^{\prime}_{ij}-w_{ij}\right)=p^{\prime}(C)-w(C).$ (12)
We therefore focus on checking if a graph $G^{\prime}=(N^{\prime},E^{\prime})$ with edge costs $c\in\mathbb{R}^{E^{\prime}}$ contains a negative cost cycle, and we introduce next a (compact) LP that achieves this goal. Recall that a cut $B\subseteq E^{\prime}$ is a set of edges of the form $\delta(X)$ for some $\emptyset\neq X\subset N^{\prime}$. We need the following theorem.
@@
$\begin{split}&x_{e}-x(B\setminus e)\leq 0,\text{ for each cut }B,\text{ for every edge }e\in B,\\ &x\geq 0.\end{split}$ (13)
Using the system of constraints from theorem 3, we can design an LP formulation as in [1, section 3], where the authors define the LP below with the goal of minimizing the cost of a cycle [^2] In [1] they also add the constraint $\sum_{e\in E^{\prime}}x_{e}=1$, because they are interested in cycles of minimum mean weight.
@@
$\begin{split}&\sum_{j:ij\in E^{\prime}\setminus\overline{e}}\left(y_{ij}-y_{ji}\right)=\begin{cases}0,&\text{ if }i\neq u,i\neq v,\\ x_{\overline{e}},&\text{ if }i=u,\\ -x_{\overline{e}},&\text{ if }i=v,\\ \end{cases}\quad\text{ for all }i\in N^{\prime},\\ &0\leq y_{ij},y_{ji}\leq x_{e},\quad\text{ for all }e=ij\in E^{\prime}\setminus\overline{e}.\end{split}$ (16)
@@
$\begin{split}\min\quad&\sum_{e\in E^{\prime}}c_{e}x_{e},\\ \text{s.t.}\quad&\sum_{j:ij\in E^{\prime}\setminus\overline{e}}\left(y_{ij}^{\overline{e}}-y_{ji}^{\overline{e}}\right)=\begin{cases}0,&\text{ if }i\neq u,i\neq v,\\ x_{\overline{e}},&\text{ if }i=u,\\ -x_{\overline{e}},&\text{ if }i=v,\\ \end{cases}\\ &\hskip 156.49014pt\text{ for all }i\in N^{\prime}\text{ and }\overline{e}=uv\in E^{\prime},\\ &0\leq y_{ij}^{\overline{e}},y_{ji}^{\overline{e}}\leq x_{e},\quad\text{ for all }e=ij\in E^{\prime}\setminus\overline{e}\text{ and }\overline{e}\in E^{\prime}.\end{split}$ (17)
@@
$\begin{gathered}p(N)=\nu(N),\\ p_{i}\geq 0,\quad\text{for all }i\in N,\\ p_{i}+p_{j}\geq w_{ij},\quad\text{for all }ij\in E,\\ \begin{drcases}&\begin{drcases}\gamma_{i}^{\overline{e}}-\gamma_{j}^{\overline{e}}-\lambda_{ij}^{\overline{e}}\\ \gamma_{j}^{\overline{e}}-\gamma_{i}^{\overline{e}}-\lambda_{ji}^{\overline{e}}\end{drcases}\leq 0,\quad\text{ for all }ij\in E^{\prime}\setminus\overline{e}\text{ and }\overline{e}\in E^{\prime},\\ &\gamma_{u}^{\overline{e}}-\gamma_{v}^{\overline{e}}+\sum_{ij\in E^{\prime}\setminus\overline{e}}\lambda_{ij}^{\overline{e}}+\lambda_{ji}^{\overline{e}}\leq(p_{u}+p_{v})/2-w_{\overline{e}},\quad\text{ for all }\overline{e}=uv\in E^{\prime},\\ &\lambda_{ij}^{\overline{e}},\lambda_{ji}^{\overline{e}}\geq 0,\quad\text{ forall }ij\in E^{\prime}\setminus\overline{e}\text{ and }\overline{e}\in E^{\prime}.\end{drcases}\\ \hfill\forall\ G^{\prime}=(N^{\prime},E^{\prime})\in\mathcal{G}.\end{gathered}$ (19)

### v2

@@ 2 Separating over the Core
Here, $\mathcal{C}$ stands for the set of cycles $C\subseteq E$ in $G$ with $b_{i}=2$ for all $i\in V(C)$, and $\mathcal{P}$ stands for the set of paths $P\subseteq E$ with $b_{i}=2$ for all inner vertices on $P$. We shortened $p(V(C))$ and $p(V(P))$ to $p(C)$ and $p(P)$, respectively. With this observation, separating over the core for a given vector $p$ reduces to checking whether $p(N)=\nu(N)$, which can be done in polynomial time (a maximum-weight $b$-matching in a graph $G$ can be computed in polynomial time, see e.g. [12]), and to separating over the set of constraints for cycles, and the set of constraints for paths.
@@
$i_{0}j^{(1)}$ for $\text{for }i_{0}j\in E$ $\text{with weight }\overline{w}_{i_{0}j}:=p_{i_{0}}+p_{j}/2-w_{i_{0}j},$
$i^{(r-1)}j^{(r)}$ for $\text{for }ij\in E$ $\text{with weight }\overline{w}_{ij}:=(p_{i}+p_{j})/2-w_{ij},$
$i^{(k-1)}j_{0}$ for $\text{for }ij_{0}\in E$ $\text{with weight }\overline{w}_{ij_{0}}:=p_{i}/2+p_{j_{0}}-w_{ij_{0}}.$
@@
$\begin{split}&p_{i}\geq 0,\quad\text{for all }i\in N,\\ &p_{i}+p_{j}\geq w_{ij},\quad\text{for all }ij\in E.\end{split}$
@@
$\begin{split}&w(C)=w_{e_{1}}+\cdots+w_{e_{k}}+w_{st}=w(P)+0=w(P),\\ &p^{\prime}(C)=p^{\prime}_{e_{1}}+\cdots+p^{\prime}_{e_{k}}+p^{\prime}_{st}=p(P),\end{split}$
@@
$\begin{split}&w(P)=w(C)-w_{st}=w(C)-0=w(C),\\ &p(P)=p^{\prime}(C),\end{split}$
@@ 3 Compact Extended Formulation
$w(C)/p^{\prime}(C)>1\iff 0>\sum_{ij\in C}c_{ij}=\sum_{ij\in C}\left(p^{\prime}_{ij}-w_{ij}\right)=p^{\prime}(C)-w(C).$
We therefore focus on checking if a graph $G^{\prime}=(N^{\prime},E^{\prime})$ with edge costs $c\in\mathbb{R}^{E^{\prime}}$ contains a negative cost cycle. There are several efficient ways to detect negative cost cycles. For example one can rely on the notion of potential in undirected graphs with general edge weights, as described in [15]. A combinatorial algorithm also follows from [4] (that actually works for a broader class of generalized matching problems). For our extended formulation, we rely on the LP formulation designed in [1] for finding negative cost cycles. Recall that a cut $B\subseteq E^{\prime}$ is a set of edges of the form $\delta(X)$ for some $\emptyset\neq X\subset N^{\prime}$. We need the following theorem.
@@
$\begin{split}&x_{e}-x(B\setminus e)\leq 0,\text{ for each cut }B,\text{ for every edge }e\in B,\\ &x\geq 0.\end{split}$
Using the system of constraints from theorem 3, we can design an LP formulation as in [1, section 3], where the authors define the LP below with the goal of minimizing the cost of a cycle. (In [1] they also add the constraint $\sum_{e\in E^{\prime}}x_{e}=1$, because they are interested in cycles of minimum mean weight, but here this constraint is not needed.)
@@
$\begin{split}&\sum_{j:ij\in E^{\prime}\setminus\overline{e}}\left(y_{ij}-y_{ji}\right)=\begin{cases}0,&\text{ if }i\neq u,i\neq v,\\ x_{\overline{e}},&\text{ if }i=u,\\ -x_{\overline{e}},&\text{ if }i=v,\\ \end{cases}\quad\text{ for all }i\in N^{\prime},\\ &0\leq y_{ij},y_{ji}\leq x_{e},\quad\text{ for all }e=ij\in E^{\prime}\setminus\overline{e}.\end{split}$
@@
$\begin{split}\min\quad&\sum_{e\in E^{\prime}}c_{e}x_{e},\\ \text{s.t.}\quad&\sum_{j:ij\in E^{\prime}\setminus\overline{e}}\left(y_{ij}^{\overline{e}}-y_{ji}^{\overline{e}}\right)=\begin{cases}0,&\text{ if }i\neq u,i\neq v,\\ x_{\overline{e}},&\text{ if }i=u,\\ -x_{\overline{e}},&\text{ if }i=v,\\ \end{cases}\\ &\hskip 156.49014pt\text{ for all }i\in N^{\prime}\text{ and }\overline{e}=uv\in E^{\prime},\\ &0\leq y_{ij}^{\overline{e}},y_{ji}^{\overline{e}}\leq x_{e},\quad\text{ for all }e=ij\in E^{\prime}\setminus\overline{e}\text{ and }\overline{e}\in E^{\prime}.\end{split}$
@@
$\begin{gathered}p(N)=\nu(N),\\ p_{i}\geq 0,\quad\text{for all }i\in N,\\ p_{i}+p_{j}\geq w_{ij},\quad\text{for all }ij\in E,\\ \begin{drcases}&\begin{drcases}\gamma_{i}^{\overline{e}}-\gamma_{j}^{\overline{e}}-\lambda_{ij}^{\overline{e}}\\ \gamma_{j}^{\overline{e}}-\gamma_{i}^{\overline{e}}-\lambda_{ji}^{\overline{e}}\end{drcases}\leq 0,\quad\text{ for all }ij\in E^{\prime}\setminus\overline{e}\text{ and }\overline{e}\in E^{\prime},\\ &\gamma_{u}^{\overline{e}}-\gamma_{v}^{\overline{e}}+\sum_{ij\in E^{\prime}\setminus\overline{e}}\lambda_{ij}^{\overline{e}}+\lambda_{ji}^{\overline{e}}\leq(p_{u}+p_{v})/2-w_{\overline{e}},\quad\text{ for all }\overline{e}=uv\in E^{\prime},\\ &\lambda_{ij}^{\overline{e}},\lambda_{ji}^{\overline{e}}\geq 0,\quad\text{ forall }ij\in E^{\prime}\setminus\overline{e}\text{ and }\overline{e}\in E^{\prime}.\end{drcases}\\ \hfill\text{ for all }\ G^{\prime}=(N^{\prime},E^{\prime})\in\mathcal{G}.\end{gathered}$ (7)